I am stuck on correct answer for:
if x - 2/7 x = 5, then x =
I believe answer is "7" but I need to verify my work. Can you show me the breakdown to get correct resolve?

Answer by checkley71(8403)

X-(2/7)X=5
X-2X/7=5
(7X-2X)/7=5
5X/7=5
5X=35
X=35/5
X=7
PROOF
7-(2/7)7=5
7-14/7=5
7-2=5
5=5